# Detecting Hidden Triggers: Mapping Non-Markov Reward Functions to Markov

### v2

@@ 6 Conclusions
In this work, we introduced a formulation for learning RMs from non-Markov rewards by solving an ILP to learn an ARMDP. The ARMDP represents the cross-product between an RM and the underlying state space of an NMRDP. Importantly, we learn the ARMDP, and thus the RM, without knowledge of $P$, by identifying hidden triggers that govern the rules of $L(S\times A\times S)\rightarrow 2^{P}$. We also demonstrated how effective it can be to learn our automaton according to $H=(S\times A\times\mathbb{R})^{*}$, rather than $H=(S\times A)^{*}$, a limitation of DFAs. We proved that the ARMDP shares an equivalent expectation to the NMRDP for $T_{o}$ and can therefore be used as a proxy for maximizing reward expectations. We validated our approach by applying ARMDPQ-Learning to train Q-Learning agents over black-box RMs. We also evaluated the representative power of the ARMDP by representing the state space of a DQN as the state space of an ARMDP. We compared the training profiles of these DQNs to the training profiles of various RDQN architectures that had to learn Markov representations over the sparse histories they encode.
@@
One important aspect of this work is that the RMs we learn are inherently more interpretable than the sparse representation of histories in $H=(S\times A\times\mathbb{R})^{*}$. This is in part because we learn minimal RMs that encode $H$ under the finite, structured, and symbolic representation of an automaton. However, more importantly, interpreting an RM doesn’t require a comprehensive understanding of $S$ and $A$ outside of what the labeler encodes. This is a consequence of the RM being entirely decomposable from the ARMDP. For example, the RM in Figure 1 Task (b) conveys that we need to bring coffee to the office without ever needing to describe where the coffee might be sourced from or where the office is. Since RL is predicated on the assumption that the reward function is the most concise, robust, and transferable definition of a task [20], we posit that RMs emerge as strong candidates for reward function representation.
@@
RM representations also carry important weight for the field of Inverse RL (IRL). In IRL, the goal is to learn a reward function from behavioral data that motivated the behavior [20, 26, 8]. While most approaches assume a Markov reward function, some of this work has been focused on learning rewards in Partially Observable MDPs (POMDPs) [8] and in Semi MDPs (SMDPs) [26]. We think that relaxing the Markov assumption in a principled way, analogous to this work, could yield insight into hidden abstractions made by a decision-maker. Moreover, given their interpretable form, it might shed light on what these hidden abstractions represent.
@@
Lastly, we generally make note that this work is limited in that it assumes $S$ and $A$ are discrete. As such, we would like to consider automata extraction methods, in the future, using RNNs [28, 21, 31] as this would extend our approach into the continuous case. Weiss et al. have demonstrated promising results in this regard by applying $L*$ as an exact learning algorithm for deep recurrent models [31].
@@ Appendix A Full ILP
$\sum_{i=1}^{K}O_{m,n,i,j}=\sum_{i=1}^{K}O_{m,n+1,j,i}$ $\forall m=1,2,...,M,\forall n=1,2,...,l_{m}-1,\forall j=1,2,...,K$ $\hskip 2.0pt(4)$
@@
The goal of our Integer Linear Program (ILP) formulation is to infer a valid Abstract Reward Markov Decision Process (ARMDP), as defined in Definition 1, from a set of observed trajectories with observed reward conflicts. Reward conflicts are a result of an agent observing rewards under the lens of $S\times A\times S$ when in reality rewards are sensitive to histories of the form $H=(S\times A\times\mathbb{R})^{*}$. We introduce some notation that our ILP leverages. We derive candidate ARMDPs by solving an ILP over a set of observed trajectory traces, $T_{o}$. Let, $\tau_{m}\in T_{o}$, be the m-th trace in $T_{o}$ of length $l_{m}$. Further, let $\tau_{m,n}=(s_{n}^{m},a_{n}^{m},r_{n+1}^{m},s_{n+1}^{m})$ be the n-th tuple of $\tau_{m}$. We define a reward conflict to be any two $\tau_{m,n}$ and $\tau_{\tilde{m},\tilde{n}}$ where $(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})=(s_{\tilde{n}}^{\tilde{m}},a_{\tilde{n}}^{\tilde{m}},s_{\tilde{n}+1}^{\tilde{m}})$, but $r_{n+1}^{m}\neq r_{\tilde{n}+1}^{\tilde{m}}$.
We define the matrix, $O=M\times N\times K\times K$, where $M=|T_{o}|$, $N$ reflects the largest $l_{m}$, $\forall\tau_{m}\in T_{o}$, and $K=|U|$ reflects the size of the finite state space for some candidate RM. $O_{m,n}$, then, is a $K\times K$ matrix that reflects the mapping of $\tau_{m,n}=(s_{n}^{m},a_{n}^{m},r_{n+1}^{m}s_{n+1}^{M})$ onto the ARMDP, where both $s_{n}^{m}$ and $s_{n+1}^{m}$ are assigned an RM state. Additionally, we make use of two sets to gather necessary variables from $T_{o}$ for the construction of our ILP. These are $(S,A,S)_{o}=\{(s_{n}^{m},a_{n}^{m},s_{n+1}^{m}),\forall\tau_{m,n}\in\tau_{m},\forall\tau_{m}\in T_{o}\}$ and $r(s,a,s^{\prime})=\{r_{n}^{m},\forall\tau_{m,n}\in\tau_{m},\forall\tau_{m}\in T_{o}\mid(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})=(s,a,s^{\prime})\}$ where $(S,A,S)_{o}$ reflects the set of all observed transition tuples in $T_{o}$ and $r(s,a,s^{\prime})$ is the set of all observed rewards for some observed transition tuple in $T_{o}$.
@@
The implication of this construction on the form of inferred ARMDP is that the underlying RM can never return to a prior state, once left. To reiterate, this does not limit the expressivity of the inferred ARMDP, it simply requires larger $U$ in some instances. Consider the RMs in Figure 5. Task (a) represents a patrol task identical to Task (e) in Figure 1 in the main paper. Even though Task (a) has atomic cycles (e.g., $u_{1}\rightarrow u_{1}$), once a state is left, the RM never returns to it. For RMs like this, even with $j\geq i$, we will still be able to represent the RM minimally. Conversely, Task (b) is a continuous patrol task, where on completion of a lap, the RM begins again in $u_{1}$. Here, because the RM reuses previous states, the ARMDPs that we produce will continuously unroll these states and expand $U$ as needed. Instead of $u_{1}\rightarrow u_{2}\rightarrow u_{3}\rightarrow u_{4}\rightarrow u_{1}...$, the RM we infer will be represented as $u_{1}\rightarrow u_{2}\rightarrow u_{3}\rightarrow u_{4}\rightarrow u_{5}...\hskip 2.0pt$. However, because $T_{o}$ represents a finite set of finite trajectories, we will always be able to learn an ARMDP from it, though perhaps at the expense of excess computation in cycle decomposition. Regardless, all inferred ARMDPs will be equally expressive.
@@ Appendix C ARMDPs as NMRDP Proxies
The goal of this section is to prove that the ARMDP is a sufficient proxy of the NMRDP for the purposes of maximizing reward expectations. Because we assume the form of $R$ in the NMRDP to be an RM, then any observed reward in $T_{o}$ would be consistently reproducible assuming identical histories. Therefore, so long as the likelihood of producing trajectories under $T$ is identical to $\tilde{T}$, reward expectations are preserved. We restate the the weighted reward sum for some $\tau_{m}\in T_{o}$:
$R(\tau_{m})=\sum_{n=1}^{l_{m}}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})$ (14)
Similarly, we restate the weighted reward sum for some $\tilde{\tau}_{m}$ derived by mapping $\tau_{m}$ onto an ARMDP to be:
$R(\tilde{\tau}_{m})=\sum_{n=1}^{l_{m}}\sum_{i=1}^{K}\sum_{j=1}^{K}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\cdot O_{m,n,i,j}$ (15)
@@
By multiplying $O_{m,n,i,j}$ into $T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})$, $O_{m,n,i,j}$ serves as a logical gate to produce the dynamics of $\tilde{T}$. With this we introduce Theorem 1:
@@
$R(\tau_{m})=R(\tilde{\tau}_{m})$ $\forall\tau_{m}\in T_{o}$ where $\tilde{\tau}_{m}$ is derived by mapping $\tau_{m}$ onto an ARMDP.
@@
$R(\tilde{\tau}_{m})$ $=\sum_{n=1}^{l_{m}}\sum_{i=1}^{K}\sum_{j=1}^{K}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\cdot O_{m,n,i,j}$
$=\sum_{n=1}^{l_{m}}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\sum_{i=1}^{K}\sum_{j=1}^{K}\cdot O_{m,n,i,j}$
@@
$=\sum_{n=1}^{l_{m}}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})=R(\tau_{m})$
@@ Appendix D Experiment 1 Specifications
For all tasks in Experiment 1, we used $L:S\rightarrow 2^{P}$ and $j\geq i$ toggle construction. Hyper parameters include $\gamma=0.95$, $\alpha=0.1$ and $\epsilon=0.1$. The only exception is for Task (b) where we use $\epsilon=0.2$ until the first conflict, to promote exploring. This is because the shortest path to the goal state includes the necessary tile to trigger the non-Markov reward. As such, the Markov representation under $S$ and $A$ appears sufficient to demonstrate optimal behavior. Even with $\epsilon=0.2$, for Task (b), 100,000 episodes was oftentimes not enough to trigger a conflict. To further accommodate this we put a sliding window on the total episode counts used to produce Figure 2 in the main paper such that the 100,000 episodes contained the first conflict. This was strictly to produce a consistent figure over 10 randomly seeded trials. We include an extended version of Figure 2 from the main paper in Figure 6 containing information on $|U|$, $|T_{o}|$, as well as Cumulative Solve Times (in seconds). Machine specifications used in the experiment: 32 core Intel(R) Xeon(R) CPU E5-2640 v3 @ 2.60 GHz with 512 GB of Memory.
Figure 6: Results of ARMDPQ-Learning evaluated on Tasks (b-e) from the Officeworld domain on 10 random trials. Episode Reward Avg column shows average reward +/- a stddev. $|U|$ column shows current assumed size of $U$. $|T_{o}|$ column shows current size of $T_{o}$. Cumulative Solve Time(s) represents the total solve time for all ARMDP inferences, in seconds.
@@ Appendix E Experiment 2 Specifications
Each model in Experiment 2 was trained using a Double Deep Q-Network (DDQN) architecture, meaning each model had a target and non-target model supporting the Deep Q-Network (DQN) architecture. All target models were trained slowly off of their respective non-target model using a smoothing factor of 0.01. Each model across the types (ADQN, DQN, RDQN, GRUDQN, and LSTMDQN) were all trained using a learning rate of 1e-3 with Adam optimizer. ADQN and DQN used a batch size of 32, whereas the RDQN, GRUDQN, and LSTMDQN models used a batch update by stacking a single trajectory. Other hyper parameters include $\gamma=0.95$ and $\epsilon=0.1$. Machine specifications used in the experiment: 32 core Intel(R) Xeon(R) CPU E5-2640 v3 @ 2.60 GHz with 512 GB of Memory.
@@ Appendix F Experiment 3 Specifications
For all tasks in Experiment 3, we used $L:S\rightarrow 2^{P}$ and $j\geq i$ toggle construction. Hyper parameters include $\gamma=0.95$, $\alpha=0.1$ and $\epsilon=0.1$. We applied a similar sliding window as we did in Experiment 1 but under a lens of 50,000 episodes. We included an extended version of Table 1 in Figure 7. Machine specifications used in the experiment: 32 core Intel(R) Xeon(R) CPU E5-2640 v3 @ 2.60 GHz with 512 GB of Memory.
Figure 7: Results of ARMDPQ-Learning evaluated on Tasks (b) and (c) from the Breakfastworld domain on 10 random trials. Task (b) and (c) Full show results when using the full RM for Task (b) and (c). Task (b) Cumulative shows results for using the RM for Task(b) where all rewards are 0 except the terminal cumulative rewards. Episode Reward Avg shows average reward +/- a stddev. $|U|$ shows current assumed size of $U$. $|T_{o}|$ shows current size of $T_{o}$. Cumulative Solve Time(s) shows the total solve time, for all ARMDPs, in seconds.

### v3

@@ 6 Conclusions
In this work, we introduced a formulation for learning RMs from non-Markov rewards by solving an ILP to learn an ARMDP. The ARMDP represents the cross-product between an RM and the underlying state space of an NMRDP. Importantly, we learn the ARMDP, and thus the RM, without knowledge of $P$ or $L$, by identifying hidden triggers that govern the rules of $L:(S\times A\times S)\rightarrow 2^{P}$. We also demonstrated how effective it can be to learn our automata according to $H=(S\times A\times\mathbb{R})^{*}$, rather than $H=(S\times A)^{*}$, a limitation of DFAs. We proved that the ARMDP is a suitable proxy for maximizing the reward expectations of the NMRDP for $T_{o}$. We validated our approach by applying ARMDPQ-Learning to train Q-Learning agents over black-box RMs. We also evaluated the representative power of the ARMDP by representing the state space of a DQN as the state space of an ARMDP. We compared the training profiles of these DQNs to the training profiles of various RDQN architectures that had to learn Markov representations over the sparse histories they encode.
@@
An important aspect of this work is that the RMs we learn are inherently more interpretable than the sparse representation of histories in $H=(S\times A\times\mathbb{R})^{*}$. This is in part because we learn minimal RMs that encode $H$ under the finite, structured, and symbolic representation of an automaton, but also because interpreting an RM doesn’t require a comprehensive understanding of $S$ and $A$ outside of what the labeler encodes. This is a consequence of the RM being entirely decomposable from the ARMDP. For example, the RM in Figure 1 Task (b) conveys the need to bring coffee to the office without ever needing to describe the sequence of states and actions to achieve that goal. Since RL is predicated on the assumption that the reward function is the most concise, robust, and transferable definition of a task [19], we posit that RMs emerge as strong candidates for reward function representation and interpretation.
@@
RM representations also carry important weight for the field of Inverse RL (IRL). In IRL, the goal is to learn a reward function from behavioral data that motivated that behavior [19, 25, 8]. While most approaches assume a Markov reward function, some of this work has been focused on learning rewards in Partially Observable MDPs (POMDPs) [8] and in Semi MDPs (SMDPs) [25]. We think that relaxing the Markov assumption in a principled way, analogous to this work, could yield insight into hidden abstractions made by a decision-maker. Moreover, given their interpretable form, it might shed light on what these hidden abstractions represent.
@@
Lastly, we generally make note that this work is limited in that it assumes $S$ and $A$ are discrete. As such, we would like to consider automata extraction methods, in the future, using recurrent models [27, 20, 30] as this would extend our approach into the continuous case. Weiss et al. have demonstrated promising results in this regard [30]. Similarly, this work assumes $R$ to be deterministic to some history. In scenarios where $R$ is stochastic or noisy, we posit that we might be able to handle such cases by sampling histories to assess if the reward distributions differ in some statistical sense.
{ack}
This research was funded in part by the Air Force Office of Scientific Research Grant No. FA9550-20-1-0032 and the Office of Naval Research Grant No. N00014-19-1-2211.
@@ Appendix A Full ILP
$\sum_{i=1}^{K}O_{m,n,i,j}=\sum_{j^{\prime}=1}^{K}O_{m,n+1,j,j^{\prime}}$ $\forall m=1,2,...,M,\forall n=1,2,...,l_{m}-1,\forall j=1,2,...,K$ $\hskip 2.0pt(4)$
@@
The goal of our Integer Linear Program (ILP) formulation is to infer a valid Abstract Reward Markov Decision Process (ARMDP), as defined in Definition 1, in the main paper, from a set of observed trajectories, $T_{o}$, with observed reward conflicts. Reward conflicts are a result of an agent observing rewards under the lens of $S\times A\times S$, when in reality, rewards are sensitive to histories of the form $H=(S\times A\times\mathbb{R})^{*}$. We introduce some notation that our ILP leverages. Let, $\tau_{m}\in T_{o}$, be the m-th trace in $T_{o}$ of length $l_{m}$. Further, let $\tau_{m,n}=(s_{n}^{m},a_{n}^{m},r_{n+1}^{m},s_{n+1}^{m})$ be the n-th tuple of $\tau_{m}$. We define a reward conflict to be any two $\tau_{m,n}$ and $\tau_{\tilde{m},\tilde{n}}$ where $(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})=(s_{\tilde{n}}^{\tilde{m}},a_{\tilde{n}}^{\tilde{m}},s_{\tilde{n}+1}^{\tilde{m}})$, but $r_{n+1}^{m}\neq r_{\tilde{n}+1}^{\tilde{m}}$.
We define the matrix, $O=M\times N\times K\times K$, where $M=|T_{o}|$, $N$ reflects the largest $l_{m}$, $\forall\tau_{m}\in T_{o}$, and $K=|U|$ reflects the size of the finite state space for some candidate Reward Machine (RM). $O_{m,n}$, then, is a $K\times K$ matrix that reflects the mapping of $\tau_{m,n}=(s_{n}^{m},a_{n}^{m},r_{n+1}^{m}s_{n+1}^{M})$ onto the ARMDP, where both $s_{n}^{m}$ and $s_{n+1}^{m}$ are assigned an RM state. Additionally, we make use of two sets to gather necessary variables from $T_{o}$ for the construction of our ILP. These are $(S,A,S)_{o}=\{(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\mid m\in\{1,2,...,M\},n\in\{1,2,...,l_{m}\}\}$ and $r(s,a,s^{\prime})=\{r_{n+1}^{m}\mid m\in\{1,2,...,M\},n\in\{1,2,...,l_{m}\},(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})=(s,a,s^{\prime})\}$ where $(S,A,S)_{o}$ reflects the set of all observed transition tuples, $(s,a,s^{\prime})\in S\times A\times S$, that exist in $T_{o}$, and $r(s,a,s^{\prime})$ is the set of all observed rewards for some given $(s,a,s^{\prime})$.
@@
The implication of this construction on the form of inferred ARMDP is that the underlying RM can never return to a prior state, once left. To reiterate our sentiments from the main paper, this does not limit the expressivity of the inferred ARMDP, it simply requires a larger $U$ in some instances. Consider the RMs in Figure 5. Task (a) represents a patrol task identical to Task (e) in Figure 1, in the main paper. Even though Task (a) has atomic cycles (e.g., $u_{1}\rightarrow u_{1}$), once a state is left, the RM never returns to it. For RMs like this, even with $j\geq i$ imposed, we still learn the minimal RM. Conversely, Task (b) is a continuous patrol task, where on completion of a single patrol cycle, the RM begins again in $u_{1}$. Here, because the RM reuses previous states, the ARMDPs that we produce will continuously unroll these states and expand $U$ as needed. Instead of $u_{1}\rightarrow u_{2}\rightarrow u_{3}\rightarrow u_{4}\rightarrow u_{1}...$, the RM we infer will be represented as $u_{1}\rightarrow u_{2}\rightarrow u_{3}\rightarrow u_{4}\rightarrow u_{5}...\hskip 2.0pt$. However, because $T_{o}$ represents a finite set of finite trajectories, we will always be able to learn an ARMDP from it, though perhaps at the expense of excess computation in cycle decomposition. Regardless, all inferred ARMDPs will be equally expressive.
@@ Appendix C ARMDPs as NMRDP Proxies
The goal of this section is to further elaborate on how the ARMDP is a sufficient proxy of the NMRDP for the purposes of maximizing reward expectations. We start by reintroducing the weighted Reward Sum (RS) for some $\tau_{m}\in T_{o}$:
$RS(\tau_{m})=\sum_{n=1}^{l_{m}}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})$ (14)
Given that we assume the form of $R$ to be an RM and that $\delta_{u}$ and $\delta_{r}$, in an RM, are deterministic functions, $RS(\tau_{m})$ reliably captures the reward observations essential for constructing a reward expectation, assuming the trajectory sequence is preserved. Similarly, we construct a weighted Abstracted Reward Sum (ARS) for some $\tau_{m}\in T_{o}$ under the representation of the ARMDP to be:
$ARS(\tau_{m})=\sum_{n=1}^{l_{m}}\sum_{i=1}^{K}\sum_{j=1}^{K}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\cdot O_{m,n,i,j}$ (15)
@@
By multiplying $O_{m,n,i,j}$ into $T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})$, $O_{m,n,i,j}$ serves as a logical gate to produce the dynamics of $\tilde{T}$. With this we reintroduce Theorem 1:
@@
$RS(\tau_{m})=ARS(\tau_{m})$ $\forall\tau_{m}\in T_{o}$.
@@
$ARS(\tau_{m})$ $=\sum_{n=1}^{l_{m}}\sum_{i=1}^{K}\sum_{j=1}^{K}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\cdot O_{m,n,i,j}$
$=\sum_{n=1}^{l_{m}}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\sum_{i=1}^{K}\sum_{j=1}^{K}O_{m,n,i,j}$
@@
$=\sum_{n=1}^{l_{m}}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})=RS(\tau_{m})$
Assuming $T_{o}$ is representative sample of the reward dynamics of $R$, in an NMRDP, by Theorem 1, the ARMDP is a suitable proxy for maximizing reward expectations for the NMRDP.
@@ Appendix D Experiment 1 Specifications
For all tasks in Experiment 1, we used $L:S\rightarrow 2^{P}$ and $j\geq i$ toggle construction. We adapted the Officeworld environment from https://github.com/RodrigoToroIcarte/reward_machines.git. We noticed that in this implementation, the * tile was a terminal state but resulted in a neutral reward of 0. We instead marked all * tiles as terminal states with a reward of -1 to deter the agent from stepping onto them. Hyper parameters included $\gamma=0.95$, $\alpha=0.1$ and $\epsilon=0.1$. The only exception is for Task (b) where we used $\epsilon=0.2$ until the first conflict, to promote exploring. This is because the shortest path to the goal state included the necessary tile to trigger the non-Markov reward. As such, the Markov representation under $S$ and $A$ appeared sufficient to demonstrate optimal behavior. Even with $\epsilon=0.2$, for Task (b), 100,000 episodes was oftentimes not enough to trigger a conflict. To further accommodate this we put a sliding window on the total episode counts used to produce Figure 2, in the main paper, such that the 100,000 episodes contained the first conflict. This was strictly to produce a consistent figure over 10 randomly seeded trials. We include an extended version of Figure 2 from the main paper in Figure 6 containing information on $|U|$, $|T_{o}|$, as well as Cumulative Solve Times (in seconds). Machine specifications used in the experiment: 32 core Intel(R) Xeon(R) CPU E5-2640 v3 @ 2.60 GHz with 512 GB of Memory.
Figure 6: Results of ARMDPQ-Learning evaluated over 10 random trials on Tasks (b-e) in the Officeworld domain. Average Rewards column is displayed with standard deviations. $|U|$ column shows current assumed size of $U$. $|T_{o}|$ column shows current size of $T_{o}$. Cumulative Solve Time(s) represents the total solve time for all ARMDP inferences, in seconds.
@@ Appendix E Experiment 2 Specifications
Each model in Experiment 2 was trained using a Double Deep Q-Network (DDQN) architecture, meaning that it leveraged a target and non-target model. All target models were trained slowly off of their respective non-target model using a smoothing factor of 0.01. Each model across the types (ADQN, DQN, RDQN, GRUDQN, and LSTMDQN) were all trained using a learning rate of 1e-3 with Adam optimizer. ADQN and DQN used a batch size of 32, whereas the RDQN, GRUDQN, and LSTMDQN models used a batch update by stacking a single trajectory. Other hyper parameters included $\gamma=0.95$ and $\epsilon=0.1$. Machine specifications used in the experiment: 32 core Intel(R) Xeon(R) CPU E5-2640 v3 @ 2.60 GHz with 512 GB of Memory.
@@ Appendix F Experiment 3 Specifications
For all tasks in Experiment 3, we used $L:S\rightarrow 2^{P}$ and $j\geq i$ toggle construction. Hyper parameters included $\gamma=0.95$, $\alpha=0.1$ and $\epsilon=0.1$. We applied a similar sliding window, as we did in Experiment 1, but under a lens of 50,000 episodes. We include an extended version of Table 1 in Figure 7. Machine specifications used in the experiment: 32 core Intel(R) Xeon(R) CPU E5-2640 v3 @ 2.60 GHz with 512 GB of Memory.
Figure 7: Results of ARMDPQ-Learning evaluated over 10 random trials on Tasks (b) and (c) from the Breakfastworld domain. Task (b) and (c) Full show results when using the full RM for Task (b) and (c). Task (b) Cumulative shows results for using the RM for Task (b) where all rewards are 0 except the terminal cumulative rewards. Average Rewards column is displayed with standard deviations. $|U|$ column shows current assumed size of $U$. $|T_{o}|$ column shows current size of $T_{o}$. Cumulative Solve Time(s) represents the total solve time for all ARMDP inferences, in seconds.
Appendix G Code
All code available at https://github.com/GregHydeDartmouth/ARMDP_RL.git